# GaussianFlowOcc: Sparse and Weakly Supervised Occupancy Estimation using Gaussian Splatting and Temporal Flow

### v2

@@ 3.4 Temporal Gaussian Splatting
Cameras in autonomous driving scenarios usually share a relatively small frustum overlap, which complicates learning of correct depth and scene geometry. To address this limitation and increase viewpoint overlap between training cameras, we incorporate Temporal Gaussian Splatting, a technique commonly employed in recent works [40, 5, 15]. During training, for a specified temporal horizon $T$, we load the camera parameters and 2D ground truth labels of the $T$ previous and $T$ subsequent frames of the sequence. We then use GS to rasterize our estimated 3D Gaussians into predicted depth and semantic maps for all temporal cameras and compute the same 2D rendering losses as for the current frame. This strategy effectively widens the supervisory signal by incorporating data from neighboring temporal views. While naïvely applying Temporal Gaussian Splatting can already significantly improve prediction performance, it overlooks the fact that many objects in the scene are in motion. This motion introduces inconsistencies in supervisory signals across frames, as objects predicted in the current frame might have moved in previous/subsequent frames, so the 2D losses give wrong signals sometimes.
@@ 3.5 Temporal Module and Gaussian Flow
The Temporal Module takes the Gaussian features output by the Gaussian Transformer as input and computes a 3D motion offset for each Gaussian relative to each temporal frame in the horizon. To achieve this, we first define a set of learnable $D$-dimensional time tokens $\Psi\in\mathbb{R}^{T\times D}$, one token for each frame in the temporal horizon $T$. For a given target time step, we create a copy of the Gaussians and append the corresponding time token to their features. These temporally encoded features are then passed through an MLP to estimate a 3D motion offset vector $\overrightarrow{v}(t)\in\mathbb{R}^{3}$:
$\overrightarrow{v}(t)=\mathrm{MLP}_{v}\left(\mathcal{G}_{f},\oplus\Psi(t)\right).$ (11)
@@ 3.6 Voxelization
The final output of our model consists solely of 3D Gaussian distributions (see Fig. 1), with no voxelization or dense operations used during the model’s architecture or training process. However, these Gaussians can be efficiently voxelized as a post-processing step to facilitate direct comparison with prior work on occupancy estimation benchmarks. For each voxel center with position $p_{i}=(x,y,z)$ on a defined voxel grid, the estimated 3D Gaussian distributions are queried and their opacity and semantic logits are accumulated at each point. Each voxel is assigned a label corresponding to the highest accumulated semantic logits. Voxels with an accumulated opacity below a defined threshold are classified as free. Further details about the voxelization are provided in the appendix.
@@ 4.1 Dataset and Metrics
We conduct experiments on the Occ3D-nuScenes dataset [52], a widely used benchmark for occupancy estimation based on the nuScenes dataset [6]. The dataset defines a voxel grid spanning $[-40m,40m]$ along the X and Y axes and $[-1m,5.4m]$ along the Z axis, with a voxel size of $0.4m$. Each voxel is assigned one of $17$ semantic labels from nuScenes Lidarseg [12] or classified as free space. Performance on this benchmark is evaluated using the Intersection over Union (IoU) score for each of the $17$ semantic categories, aggregated to the mean IoU (mIoU). Additionally, following recent trends in the literature, we also report the RayIoU metric [31], a ray-based metric focused on rewarding good 3D completion while penalizing thick surface predictions. Finally, we provide information on inference speed (FPS) on a single A100 GPU.
@@ 4.2 Implementation Details
We utilize a ResNet50 [17] backbone for feature extraction with an image resolution of $256\times 704$ and set the number of Gaussians to $N=10000$. The Gaussian Transformer consists of three blocks, each with induced attention layers employing $M=500$ inducing points. The Gaussian Heads are implemented as a single linear layer, while all layers in the model share a latent dimension of $256$. The Temporal Module consists of three linear layers and operates with a temporal horizon $T=6$ during training. The model is trained for $18$ epochs using four A100 GPUs.
@@
SelfOcc [20] ResNet50 OpenSeeD 10.54 45.01 - - - - 1.15
OccNeRF [67] ResNet101 GroundedSAM 10.81 22.81 - - - - 1.27
GaussianOcc [15] ResNet101 GroundedSAM 11.26 - 11.85 8.69 11.90 14.95 5.57
GaussTR [24] 2×2$\times$2 ×ViT CLIP + Metric3D + GroundedSAM 13.26 45.19 - - - - 0.20
GaussianFlowOcc (Ours) ResNet50 GroundedSAM + Metric3D 16.02 46.91 16.47 11.81 16.58 20.98 10.2
We evaluate our proposed method against state-of-the-art approaches for weakly supervised 3D occupancy prediction on the Occ3D-nuScenes dataset [52], with the results summarized in Tab. 1. Each of the compared methods relies on a pretrained semantic segmentation model to generate 2D semantic labels for training. Our approach significantly outperforms all existing models, achieving at least a $42\%$ mIoU improvement over methods that rely on RGB reconstruction and pseudo semantic labels. Notably, we surpass the recent GaussTR, which also uses Metric3D for depth supervision, by $21\%$ mIoU, improving from $13.26$ to $16.05$. Furthermore, our model achieves this superior performance without relying on expensive backbones such as CLIP and Metric3D, which GaussTR incorporates at the cost of significantly higher inference time. A similar trend is observed in the RayIoU metric, where our method improves performance by approximately $39\%$ compared to GaussianOcc.
Beyond accuracy, our model also offers a substantial efficiency advantage. Due to its lightweight architecture, it achieves inference speeds up to almost 9×9$\times$9 × faster than previous voxel-based methods. As mentioned above, especially GaussTR, remains computationally expensive due to its reliance on multiple ViT-L backbones; GaussianFlowOcc is 50×50$\times$50 × faster.
We attribute these performance gains primarily to two key innovations: Firstly, our Temporal Module significantly improves temporal consistency, yielding large performance increases as we show in Sec. 4.4.3. Secondly, the Gaussian Transformer allows for a much higher number of Gaussians to be processed efficiently, due to approximately linearly scaling induced attention modules.
A detailed per-class performance breakdown is provided in the appendix. Additionally, qualitative results presented in Fig. 1 and in the appendix highlight our model’s strong 3D scene completion abilities. In particular, our approach excels at representing small, thin, and flat objects (e.g., traffic signs, poles, and pedestrians), a major advantage of using 3D Gaussians, which can model arbitrary geometries. In contrast, voxel-based methods are inherently constrained by a grid resolution, limiting their ability to capture fine details.
@@ 4.4 Ablation Study
Induced Self-Attention Induced Temporal Attention mIoU
@@ 4.4.3 Temporal Module
We demonstrate the effectiveness of the proposed Temporal Module by training a model without it while keeping all other components unchanged. As visible in Tab. 4, using the Temporal Module leads to a relative improvement in mIoU of $13\%$ and in RayIoU of $14\%$, which emphasizes the importance of object motion compensation during training. A demonstration of rasterized views of temporally adjacent frames with and without applied estimated flow is provided in the appendix and in Fig. 1.
@@ 4.4.4 Gaussian Parameters
In contrast to previous works [15, 48], we find that allowing the model to freely adjust all Gaussian parameters yields the best results. To demonstrate this, we conduct an experiment in which we disable the Gaussian Heads for certain properties and fix their values for GS, preventing them from being learned. We fix the opacity with $1$, the scale with $0.3$ and the rotations with unit quaternions. When only estimating the mean and fixing all other properties, only a performance of $9.48$ mIoU can be achieved, showing the importance of the other properties. When allowing the model to influence the Gaussians opacity, a large improvement to $13.12$ mIoU is reached, outperforming most previous methods. Adding the other properties further increases performance, with the maximum score achieved when using all properties. As mentioned above, the positive effect of using rotation and scale is especially noticeable for thin or flat structures.
@@ 4.4.5 Time Horizon
We ablate the Temporal Gaussian Splatting horizon $T$ in the range of $\{0,2,...,10\}$ to show how temporal rendering leads to better supervision. With horizons greater than 8, the model diverges, which is likely due to large discrepancy between the time steps, making it increasingly harder to estimate (especially non-linear) motion. Additionally, some camera poses are just too distant from the current time step.
Table 5: Ablation on different horizons $T$. The best performance is achieved with $T=6$.
@@ 4.4.6 Pseudo Depth Labels
To analyze the impact of pseudo depth labels generated by Metric3D, we train a model using only pseudo semantics. The results of this experiment are presented in Tab. 6. Even without pseudo depth supervision, our model outperforms all prior works not using depth, demonstrating its strong capability. Notably, other approaches [67, 20] incorporate additional photometric losses to enhance geometric consistency, which we did not find to be beneficial. Interestingly, GaussTR [24] fails to converge without depth labels, further highlighting the robustness of our approach. This suggests that Temporal Gaussian Splatting alone provides sufficient 3D cues for geometry estimation by leveraging semantic maps from multiple cameras over an extended time horizon.
@@ 5 Conclusion and Future Work
In this work, we introduced a novel, linearly scaling Transformer architecture for efficiently modeling driving scenes using 3D Gaussian distributions, trained with Temporal Gaussian Splatting and explicit object motion modeling. Our approach significantly outperforms previous weakly supervised methods on the Occ3D-nuScenes benchmark while maintaining substantially faster inference speeds. Despite these advances, there are still areas for improvement. We found that self-supervised learning (SSL) with RGB rendering did not provide benefits, suggesting that further investigation is needed in this direction. Moreover, a more sophisticated approach to dynamic object modeling could further improve temporal consistency. Incorporating physically informed motion priors, leveraging temporal and Gaussian dependencies as well as adapting more properties than just the mean to better represent deformable objects could enhance temporal modeling. For future work, the sparse query mechanism and efficient attention proposed in this work could be applied to other downstream tasks, such as object detection, lane segmentation, and planning. This opens up exciting possibilities for using our model as a generic world representation in autonomous driving and robotics applications.
Appendix
Appendix A Detailed Quantitative Results
For completeness, we provide per-class IoU scores for all evaluated methods on the Occ3D-nuScenes dataset in Tab. A.1. Our proposed GaussianFlowOcc consistently outperforms all prior methods trained with pseudo labels, both with and without depth supervision. The most significant performance gains can be observed in small object classes, such as motorcycles and traffic cones, but also in categories. Notably, our method achieves the best or second best score across nearly all categories. The only exception are the the LiDAR-based labels, where one method surpasses our performance by a small margin. These results further highlight the effectiveness of our Gaussian-based scene representation and temporal modeling, particularly in handling fine-grained object details that are often challenging for voxel-based approaches.
@@ Appendix A Detailed Quantitative Results
GaussianFlow (Ours) C+D 16.02 - 7.23 9.33 17.55 17.94 4.50 9.32 8.51 10.66 2.0 11.80 63.89 - 31.11 35.12 14.64 12.59
@@ B.1 Predicted Occupancy
In Fig. B.1, we present qualitative examples of GaussianFlowOcc’s predicted occupancy compared to the ground truth, visualized from a third-person perspective. These examples demonstrate our model’s ability to precisely reconstruct the 3D scene geometry using 3D Gaussians, despite the absence of explicit 3D supervision. The Gaussians align naturally to capture the shapes and details of objects, effectively representing scene structures without requiring per-scene optimization. This highlights the flexibility and efficiency of our approach in modeling complex environments. Additionally, we provide a set of videos in our repository showcasing inference results on full validation scenes:

### v3

@@ 3.5 Temporal Gaussian Splatting
Cameras in autonomous driving scenarios usually share a relatively small frustum overlap, which complicates learning of correct depth and scene geometry. To address this limitation and increase viewpoint overlap between training cameras, we incorporate Temporal Gaussian Splatting, a technique commonly employed in recent works [41, 5, 15]. During training, for a specified temporal horizon $T$, we load the camera parameters and 2D ground truth labels of the $T$ previous and $T$ subsequent frames of the sequence. We then use GS to rasterize our estimated 3D Gaussians into predicted depth and semantic maps for all temporal cameras and compute the same 2D rendering losses as for the current frame. This strategy effectively widens the supervisory signal by incorporating data from neighboring temporal views. While naïvely applying Temporal Gaussian Splatting can already significantly improve prediction performance, it overlooks the fact that many objects in the scene are in motion. This motion introduces inconsistencies in supervisory signals across frames, as objects predicted in the current frame might have moved in previous/subsequent frames, so the 2D losses can induce erroneous signals.
@@ 3.6 Temporal Module and Gaussian Flow
The Temporal Module takes the Gaussian features output by the Gaussian Transformer as input and computes a 3D motion offset for each Gaussian relative to each temporal frame in the horizon. To achieve this, we first define a set of learnable $D$-dimensional time tokens $\Psi\in\mathbb{R}^{2T\times D}$, one token for each frame in the temporal horizon ([T,T][-T,T]). For a given target time step, we create a copy of the Gaussians and append the corresponding time token to their features. These temporally encoded features are then passed through an MLP to estimate a 3D motion offset vector $\overrightarrow{v}(t)\in\mathbb{R}^{3}$:
$\overrightarrow{v}(t)=\mathrm{MLP}_{v}\left(\mathcal{G}_{f}\oplus\Psi(t)\right).$ (11)
@@ 3.7 Voxelization
The final output of our model consists solely of 3D Gaussian distributions (see Fig. 1), with no voxelization or dense operations used during the model’s architecture or training process. However, these Gaussians can be efficiently voxelized as a post-processing step to facilitate direct comparison with prior work on occupancy estimation benchmarks. For each voxel center with position $p_{i}=(x,y,z)$ on a defined voxel grid, the estimated 3D Gaussian distributions are queried and their opacity and semantic logits are accumulated at each point. Each voxel is assigned a label corresponding to the highest accumulated semantic logits. Voxels with an accumulated opacity below a defined threshold are classified as free. Further details about the voxelization are provided in the supplementary material.
@@ 4.1 Dataset and Metrics
We conduct experiments on the Occ3D-nuScenes dataset [53], a widely used benchmark for occupancy estimation based on the nuScenes dataset [6]. The dataset defines a voxel grid spanning [40m,40m][-40m,40m] along the X and Y axes and [1m,5.4m][-1m,5.4m] along the Z axis, with a voxel size of $0.4m$. Each voxel is assigned one of $17$ semantic labels from nuScenes Lidarseg [12] or classified as free space. Performance on this benchmark is evaluated using the Intersection over Union (IoU) score for each of the $17$ semantic categories, aggregated to the mean IoU (mIoU). Additionally, following recent trends in the literature, we also report the RayIoU metric [31], a ray-based metric focused on rewarding good 3D completion while penalizing thick surface predictions. Finally, we provide information on inference speed (FPS) on a single A100 GPU.
@@ 4.2 Implementation Details
We utilize a ResNet-50 [17] backbone for feature extraction with an image resolution of $256\times 704$ and set the number of Gaussians to $N=10000$. The Gaussian Transformer consists of three blocks, each with induced attention layers employing $M=500$ inducing points. The Gaussian Heads are implemented as a single linear layer, while all layers in the model share a latent dimension of $256$. The Temporal Module consists of three linear layers and operates with a temporal horizon $T=6$ during training. The model is trained for $18$ epochs using four A100 GPUs. The source code is available at https://github.com/boschresearch/GaussianFlowOcc.
@@
SelfOcc [20] ResNet-50 OpenSeeD 10.54 45.01 - - - - 1.15
OccNeRF [70] ResNet-101 GroundedSAM 10.81 22.81 - - - - 1.27
GaussianOcc [15] ResNet-101 GroundedSAM 11.26 - 11.85 8.69 11.90 14.95 5.57
GaussTR [24] $2\timesViT$ CLIP + Metric3D + GroundedSAM 13.26 45.19 - - - - 0.20
GaussianFlowOcc (Ours) ResNet-50 GroundedSAM + Metric3D 17.08 46.91 16.47 11.81 16.58 20.98 10.2
We evaluate our proposed method against state-of-the-art approaches for weakly supervised 3D occupancy prediction on the Occ3D-nuScenes dataset [53], with the results summarized in Tab. 1. Each of the compared methods relies on a pretrained semantic segmentation model to generate 2D semantic labels for training. Our approach significantly outperforms all existing models, achieving at least a $51\%$ mIoU improvement over methods that rely on RGB reconstruction and pseudo semantic labels. Notably, we surpass the recent GaussTR, which also uses Metric3D for depth supervision, by $29\%$ mIoU, improving from $13.26$ to $17.08$. Furthermore, our model achieves this superior performance with a rather lightweight ResNet-50 backbone, while GaussTR uses CLIP and Metric3D as backbones during inference, which is significantly more costly. A similar trend is observed in the RayIoU metric, where our method improves performance by approximately $39\%$ compared to GaussianOcc. Beyond accuracy, our model also offers a substantial efficiency advantage. Due to its lightweight architecture, it achieves inference speeds up to almost $9\times$ faster than previous voxel-based methods. As mentioned above, especially GaussTR remains computationally expensive due to its reliance on multiple ViT-L backbones; GaussianFlowOcc is $50\times$ faster.
We attribute these performance gains primarily to two key innovations: Firstly, our Temporal Module significantly improves temporal consistency, yielding large performance increases as we show in Sec. 4.4.3. Secondly, the Gaussian Transformer allows for a much higher number of Gaussians to be processed efficiently, due to approximately linearly scaling induced attention modules. A detailed per-class performance breakdown is provided in the supplement. Additionally, qualitative results presented in Fig. 1 and in the supplement highlight our model’s strong 3D scene completion abilities. In particular, our approach excels at representing small, thin, and flat objects (e.g., traffic signs, poles, pedestrians), a major advantage of using 3D Gaussians. In contrast, voxel-based methods are inherently constrained by a grid resolution, limiting their ability to capture fine details.
@@ 4.4 Ablation Study
Induced
Self-Attention
Induced
Temporal Attention
mIoU
@@ 4.4.3 Temporal Module
We demonstrate the effectiveness of the proposed Temporal Module by training a model without it while keeping all other components unchanged. As visible in Tab. 4, using the Temporal Module leads to a relative improvement in mIoU of $20\%$ and in RayIoU of $14\%$, which emphasizes the importance of object motion compensation during training. A demonstration of rasterized views of temporally adjacent frames with and without applied estimated flow is provided in the supplement and in Fig. 1.
@@ 4.4.4 Gaussian Parameters
In contrast to previous works [15, 49], we find that allowing the model to freely adjust all Gaussian parameters yields the best results. To demonstrate this, we conduct an experiment (Tab. 4) where we disable the Gaussian Heads for certain properties and fix their values, preventing them from being learned. We set the opacity to $1$, the scale to $0.3$ and the rotations to unit quaternions. When only estimating the mean and fixing all other properties, only a performance of $9.48$ mIoU can be achieved, showing the importance of the other properties. When allowing the model to influence the Gaussians opacity, a large improvement to $13.12$ mIoU is reached, outperforming most previous methods. Adding the other properties further increases performance, with the maximum score achieved when using all properties. As mentioned above, the positive effect of using rotation and scale is especially noticeable for thin or flat structures.
@@ 4.4.5 Time Horizon
We ablate the Temporal Gaussian Splatting horizon $T$ in the range of $\{0,2,...,10\}$ to demonstrate how temporal rendering leads to better supervision, shown in Tab. 5. With horizons greater than 8, the model diverges, which is likely due to large discrepancy between the time steps, making it increasingly harder to estimate (especially non-linear) motion. Additionally, some camera poses are just too distant from the current time step.
Table 5: Ablation on different horizons $T$. The best performance is achieved with $T=6$. The model diverges with $T\geq 10$.
@@ 4.4.6 Pseudo Depth Labels
To analyze the impact of pseudo depth labels generated by Metric3D, we train a model using only pseudo semantics. The results of this experiment are presented in Tab. 6. Even without pseudo depth supervision, our model outperforms all prior works, demonstrating its strong capability. Notably, other approaches [70, 20] incorporate additional photometric losses to enhance geometric consistency, which we did not find to be beneficial. Interestingly, GaussTR [24] fails to converge without depth labels, further highlighting the robustness of our approach. This suggests that Temporal Gaussian Splatting alone provides sufficient 3D cues for geometry estimation by leveraging semantic maps from multiple cameras over an extended time horizon.
@@ 5 Conclusion and Future Work
In this work, we introduced a novel, linearly scaling Transformer architecture for efficiently modeling driving scenes using 3D Gaussian distributions, trained with Temporal Gaussian Splatting and explicit object motion modeling. Our approach significantly outperforms previous weakly supervised methods on the Occ3D-nuScenes benchmark while maintaining substantially faster inference speeds. Despite these advances, there are still areas for improvement. We found that self-supervised learning (SSL) with RGB rendering did not provide benefits, suggesting that further investigation is needed in this direction. Moreover, a more sophisticated approach to dynamic object modeling could further improve temporal consistency. Incorporating physically informed motion priors, leveraging temporal and Gaussian dependencies as well as adapting more properties than just the mean to better represent deformable objects could enhance temporal modeling. For future work, the sparse and efficient attention mechanism proposed in this work could be applied to other downstream tasks, such as object detection, lane segmentation, and planning. This opens up exciting possibilities for using our model in autonomous driving and robotics applications.
Supplementary Material
Appendix A Additional Quantitative Results
A.1 Per-Class IoU
For completeness, we provide per-class IoU scores for all evaluated methods on the Occ3D-nuScenes dataset in Tab. A.1. Our proposed GaussianFlowOcc consistently outperforms all prior methods trained with pseudo labels, both with and without depth supervision. The most significant performance gains can be observed in small object classes, such as motorcycles and traffic cones, but also in other categories. Notably, our method achieves the best or second best score across nearly all categories. These results further highlight the effectiveness of our Gaussian-based scene representation and temporal modeling, particularly in handling fine-grained object details that are often challenging for voxel-based approaches.
@@ A.1 Per-Class IoU
GaussianFlow (Ours) C+D 17.08 - 7.23 9.33 17.55 17.94 4.50 9.32 8.51 10.66 2.0 11.80 63.89 - 31.11 35.12 14.64 12.59
A.2 Depth Estimation
To further demonstrate the 3D geometric capabilities of GaussianFlowOcc, we report depth estimation results on the nuScenes dataset in Tab. A.2. Our method achieves competitive performance compared to other state-of-the-art self-supervised and weakly supervised approaches on this task. We observe that methods utilizing photometric losses [20, 70, 15] tend to produce slightly more accurate camera depth predictions than our approach. However, it is important to note that while our method may yield marginally lower depth estimation accuracy, it substantially outperforms these methods in 3D occupancy estimation, underscoring the strength of our model in comprehensive 3D scene understanding.
Table A.2: Depth estimation results on nuScenes. The best result is highlighted in bold, second best in italics.
Method Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ RMSE log $\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$
SurroundDepth [59] 0.280 4.401 7.467 0.364 0.661 0.844 0.917
SimpleOcc [14] 0.224 3.383 7.165 0.333 0.753 0.877 0.930
OccNeRF [70] 0.202 2.883 6.697 0.319 0.768 0.882 0.931
SelfOcc [20] 0.215 2.743 6.706 0.316 0.753 0.875 0.932
GaussianOcc [70] 0.197 1.846 6.733 0.312 0.746 0.873 0.931
GaussianFlowOcc (Ours) 0.278 2.522 5.232 0.389 0.677 0.826 0.898
A.3 Training with 3D Voxel Labels
Our main contributions, namely the Gaussian representation and the dynamic object compensation during temporal Gaussian Splatting, specifically target the weakly supervised setting for occupancy estimation. Nonetheless, our model can be trained using 3D voxel labels similar to previous methods [30, 29, 72, 53]. Specifically, we apply a differentiable voxelization to the final Gaussian predictions (see Sec. 3.7), enabling the computation of a cross-entropy loss with ground-truth voxel labels for supervised training. Note that for this experiment, the Temporal Module and Gaussian Splatting supervision are not used. In Tab. A.3, we present a comparison on the Occ3D-nuScenes dataset between our method and established fully supervised approaches, following the recommended protocol of not using the camera visibility mask [53]. As shown, our model outperforms prior methods such as BEVFormer [29] and OccFormer [72], while achieving competitive performance with larger models specifically designed for full supervision, including FB-Occ [30] and CTF-Occ [53].
Table A.3: Occupancy estimation performance on the Occ3D-nuScenes validation set using 3D voxel labels for training. The best result per column is highlighted in bold, second best in italics.
Method
Backbone
mIoU
others
barrier
bicycle
bus
car
cons. vehicle
motorcycle
pedestrian
traffic cone
trailer
truck
driv. surf.
other flat
sidewalk
terrain
manmade
vegetation
BEVFormer [29] ResNet-101 23.67 5.0 38.8 10.0 34.4 41.1 13.2 16.5 18.2 17.8 18.7 27.7 49.0 27.7 29.1 25.4 15.4 14.5
OccFormer [72] ResNet-101 21.93 5.9 30.3 12.3 34.4 39.2 14.4 16.4 17.2 9.3 13.9 26.4 51.0 31.0 34.7 22.7 6.8 7.0
FB-Occ [30] ResNet-50 27.09 0.0 40.9 21.2 39.2 40.8 20.6 23.8 23.6 25.0 16.6 26.4 59.4 27.6 31.4 29.0 16.7 18.4
CTF-Occ [53] ResNet-101 28.53 8.1 39.3 20.6 38.3 42.2 16.9 24.5 22.7 21.0 23.0 31.1 53.3 33.8 38.0 33.2 20.8 18.0
GaussianFlowOcc (Ours) ResNet-50 25.28 7.6 25.4 13.4 26.0 25.3 15.3 14.2 12.7 10.7 18.2 20.9 76.2 37.4 48.2 47.8 17.3 13.2
@@ B.1 Predicted Occupancy
In Fig. B.1, we present qualitative examples of GaussianFlowOcc’s predicted occupancy compared to the ground truth, visualized from a third-person perspective. These examples demonstrate our model’s ability to precisely reconstruct the 3D scene geometry using 3D Gaussians, despite the absence of explicit 3D supervision. The Gaussians align naturally to capture the shapes and details of objects, effectively representing scene structures without requiring per-scene optimization. This highlights the flexibility and efficiency of our approach in modeling complex environments. Additionally, we provide a set of videos in our GitHub repository (see Appendix D) showcasing inference results on full validation scenes: